A 7.31-nc charge is located 1.73 m from a 4.54-nC point charge. (a) Find the magnitude of the electrostatic force that one charge exerts on the other
Answer #1

here,

charge 1 , q1 = 7.31 nC = 7.31 * 10^-9 C

charge 2 , q2 = 4.54 nC = 4.54 * 10^-9 C

separation between the charges , r = 1.73 m

the magnitude of electrostatic force that one charge exerts on the other , F = K * q1 * q2 /r^2

F = 9 * 10^9 * 7.31 * 10^-9 * 4.54 * 10^-9 /1.73^2

F = 9.98 * 10^-8 N

the magnitude of electrostatic force that one charge exerts on the other is 9.98 * 10^-8 N